### v1

# Erase then Rectify: A Training-Free Parameter Editing Approach for Cost-Effective Graph Unlearning

Abstract
Graph unlearning, which aims to eliminate the influence of specific nodes, edges, or attributes from a trained Graph Neural Network (GNN), is essential in applications where privacy, bias, or data obsolescence is a concern. However, existing graph unlearning techniques often necessitate additional training on the remaining data, leading to significant computational costs, particularly with large-scale graphs. To address these challenges, we propose a two-stage training-free approach, Erase then Rectify (ETR), designed for efficient and scalable graph unlearning while preserving the model utility. Specifically, we first build a theoretical foundation showing that masking parameters critical for unlearned samples enables effective unlearning. Building on this insight, the Erase stage strategically edits model parameters to eliminate the impact of unlearned samples and their propagated influence on intercorrelated nodes. To further ensure the GNN’s utility, the Rectify stage devises a gradient approximation method to estimate the model’s gradient on the remaining dataset, which is then used to enhance model performance. Overall, ETR achieves graph unlearning without additional training or full training data access, significantly reducing computational overhead and preserving data privacy. Extensive experiments on seven public datasets demonstrate the consistent superiority of ETR in model utility, unlearning efficiency, and unlearning effectiveness, establishing it as a promising solution for real-world graph unlearning challenges.
Introduction
Machine unlearning has garnered considerable attention in recent years due to its ability to remove the influence of specific subsets of data from a well-trained machine learning model (Nguyen et al. 2022). This process is particularly important for forgetting sensitive, mislabeled, or obsoleted information. Typically, the subset of data to be unlearned is much smaller than the entire training dataset. As a result, one primary goal of machine unlearning is to efficiently eliminate the impact of these unlearned samples while preserving the model’s predictive power, thereby avoiding the costly process of retraining the model from scratch on the remaining data (Xu et al. 2024).
In the graph domain, there is also a strong demand to forget specific information. For example, users in a social network may request the removal of their personal data. In particular, Graph Neural Networks (GNNs) has become a common routine for processing and analyzing graph-structured data, leveraging message propagation and neighborhood aggregation mechanisms (Zhou et al. 2020). However, the interconnected nature of graphs determines that unlearned samples can significantly impact their neighbors through message propagation, making conventional machine unlearning methods unapplicable (Cheng et al. 2023). Consequently, graph unlearning has emerged as a specialized machine unlearning task with the goal of removing unlearned samples and mitigating their influence on other intercorrelated nodes (Said et al. 2023).
Recently, several approaches have been proposed to address the challenges of graph unlearning. For instance, GIF (Wu et al. 2023a) and CGU (Chien, Pan, and Milenkovic 2023) directly approximate the parameter changes induced by data removal and update the affected parameters. Such methods primarily focus on forgetting unlearned samples while overlooking the predictive capability on remaining data, which can significantly compromise model utility (Li et al. 2024). To resolve this problem, some recent studies (Li et al. 2024; Wang, Huai, and Wang 2023) have simultaneously considered both predictive and unlearning objectives when an unlearning request is received. However, they typically necessitate additional training on the remaining data, which leads to significant computational overhead, particularly with large graphs, even if only a small number of unlearned samples are involved. The substantial computational resources required constrain the efficiency and scalability of these methods, thereby limiting their practicality in real-world applications.
In this work, we focus on developing an efficient and scalable graph unlearning method that preserves the utility of GNNs. However, achieving this goal is non-trivial. First, the interconnected nature of graph data means that unlearned samples can significantly affect their neighbors through message propagation, making it challenging to efficiently remove both the unlearned samples and their influence on other nodes. Second, the remaining dataset after unlearning is typically much larger than the unlearned subset, which can lead to substantial computational overhead when accessing the remaining data. Balancing the need to maintain model utility with the demand for low computational overhead presents another significant challenge.
To tackle the above challenges, we propose Erase then Rectify (ETR), a training-free, two-stage approach for cost-effective graph unlearning. Specifically, we first establish a theoretical foundation demonstrating that masking parameters critical to unlearned samples enables effective graph unlearning. Building on this theory, in the Erase stage, we propose a neighborhood-aware parameter editing strategy to effectively eliminate the impact of unlearned samples and their cascading effects on the entire graph, while minimizing the impact on model utility. In the Rectify stage, we introduce a subgraph-based gradient approximation method to efficiently estimate the gradient of the unlearned model on the remaining data, which is then used to further enhance the model performance. Notably, ETR achieves graph unlearning without requiring explicit model training or access to the entire training dataset, which not only avoids substantial computational overhead but also ensures data privacy and security. Overall, the main contributions are as follows:
  • We theoretically prove that masking parameters crucial for unlearned samples enable effective graph unlearning.
  • We propose a training-free neighborhood-aware parameter editing method for graph unlearning that effectively forgets unlearned samples and their influence on intercorrelated samples.
  • We propose a gradient approximation method to enhance GNN performance on the remaining graph without requiring access to the entire training dataset, thereby ensuring efficiency and scalability on large-scale graphs.
  • Extensive experiments demonstrate the model utility, unlearning efficiency, and unlearning efficacy of ETR. Notably, ETR achieves an average of 4583.9x less time and 4.2x less memory usage than retraining from scratch.
Related work
Graph Unlearning
Graph unlearning methods can be categorized into exact and approximate approaches, aiming to obtain an unlearned model that is exactly or approximately equivalent to retraining from scratch (Said et al. 2023). Current efforts primarily focus on approximate graph unlearning. For example, GraphEraser (Chen et al. 2022) and GUIDE (Wang, Huai, and Wang 2023) follow the SISA (Bourtoule et al. 2021) paradigm, proposing graph partitioning algorithms to divide the training set into multiple shards, with a separate model trained for each shard. Besides, some approaches (Wu et al. 2023a, b; Chen et al. 2023; Chien, Pan, and Milenkovic 2023) leverage influence functions to approximate the impact of unlearning requests on model parameters. GNNDelete (Cheng et al. 2023) freezes model parameters while training additional parameters to achieve graph unlearning, and MEGU (Li et al. 2024) simultaneously trains the model’s prediction and unlearning objectives. Despite the effectiveness, these methods still face limitations when scaling to large graphs.
Training-Free Machine Unlearning
Several training-free methods have been proposed for efficient machine unlearning. For example, Fisher Forgetting (Golatkar, Achille, and Soatto 2020a) and NTK (Golatkar, Achille, and Soatto 2020b) are weight scrubbing methods that add noise to parameters informative for unlearned samples. However, Fisher Forgetting is computationally expensive and sacrifices the model’s predictive performance (Tarun et al. 2023), while NTK relies on additional models to achieve unlearning. The most relevant work to ours is SSD (Foster, Schoepf, and Brintrup 2024), which achieves unlearning by dampening parameters crucial for the unlearned dataset but not for the remaining dataset. Unlike SSD, we account for the impact of unlearned samples on other samples through message propagation. Additionally, we propose adaptively selecting hyperparameters, enhancing the method’s versatility across different datasets. Furthermore, we introduce the Rectify strategy to improve the model performance on the remaining dataset.
Preliminaries
Notations and Background
In this paper, we employ $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\mathcal{X}\right)$ to denote a graph comprising $\left|\mathcal{V}\right|$ nodes and $\left|\mathcal{E}\right|$ edges. Each node $\mathbf{v}_{i}\in\mathcal{V}$ has a $d$-dimensional feature vector $\mathbf{x}_{i}\in\mathcal{X}$. A major category of GNNs is message-passing neural networks (MPNNs) (Gilmer et al. 2017), including models such as GCN (Kipf and Welling 2017) and GAT (Velickovic et al. 2018). MPNNs follow the “propagate-transform” paradigm: in each layer of the network, they propagate and aggregate features from neighboring nodes, and then transform these aggregated features to update each node’s representation (Gilmer et al. 2017). Following previous works (Chen et al. 2022; Said et al. 2023), we focus on the task of node classification in this study. We denote the training dataset as $D$, the unlearned dataset as $D_{f}$, the remaining dataset as $D_{r}$ (i.e., $D\setminus D_{f}$), and the $k$-hop neighborhood of $D_{f}$ as $D_{k}$.
Figure 1: The framework of ETR: Forgetting the influence of unlearned samples through Erase, followed by enhancing the model performance on the remaining dataset via Rectify.
Fisher Information Matrix (FIM)
The FIM utilizes the second-order derivative of the loss function to quantify the sensitivity of each parameter, thereby providing a measure of its importance with respect to the input samples (Guo et al. 2020). Specifically, for a distribution $p(y|x,w)$, the FIM and its first-order derivative property (Kay 1993) are given as follows:
$F_{D}=$ $\mathbb{E}_{x,y}\left[-\nabla_{w}^{2}\log p\left(y|x,w\right)\right]$ (1)
$=$ $\mathbb{E}_{x,y}\left[\nabla_{w}\log p\left(y|x,w\right)\nabla_{w}\log p\left(y|x,w\right)^{T}\right]$ (2)
Notably, obtaining the FIM is computationally expensive. A common approach is to use its diagonal $diag(F)$ (Kirkpatrick et al. 2017), which can be computed efficiently using the first-order derivative. The $i$-th element of the diagonal of the FIM calculated over the dataset $D$ is denoted as $F_{D,ii}$, representing the importance of the $i$-th parameter with respect to the dataset $D$.
Problem Statement
Given a graph $\mathcal{G}$, the optimal GNN model $f_{\mathcal{G}}$ trained on $\mathcal{G}$, and a graph unlearning request $\Delta\mathcal{G}=\left(\Delta\mathcal{V},\Delta\mathcal{E},\Delta\mathcal{X}\right)$, the objective of graph unlearning is to derive a new GNN model $\widehat{f}$ that minimizes the discrepancy between $\widehat{f}$ and $f_{\mathcal{G}\setminus\Delta\mathcal{G}}$. Here, $f_{\mathcal{G}\setminus\Delta\mathcal{G}}$ represents the GNN model retrained from scratch on the remaining graph $\mathcal{G}\setminus\Delta\mathcal{G}$.
Graph unlearning can be categorized into the following three tasks: Node Unlearning: $\Delta\mathcal{G}=\left(\Delta\mathcal{V},\Delta\mathcal{E},\Delta\mathcal{X}\right)$, where $\Delta\mathcal{V}$ represents the unlearned nodes, $\Delta\mathcal{E}$ denotes the edges connecting to $\Delta\mathcal{V}$, and $\Delta\mathcal{X}$ represents the features of nodes in $\Delta\mathcal{V}$. Edge Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\Delta\mathcal{E},\emptyset\right)$, where $\Delta\mathcal{E}$ represents the set of edges to be unlearned. Feature Unlearning: $\Delta\mathcal{G}=\left(\emptyset,\emptyset,\Delta\mathcal{X}\right)$, where $\Delta\mathcal{X}$ represents the unlearned features, and will be padded with zeros. We will use the node unlearning task to illustrate the proposed method, and describe how the proposed method applies to edge unlearning and feature unlearning tasks in the Appendix.
Methodology
Framework Overview
The overall framework of ETR is depicted in Figure 1. In the Erase stage, we evaluate parameter importance with respect to input samples based on the FIM. Subsequently, we identify critical parameters concerning the unlearned samples and their influence on other samples. We then modify these parameters to erase the effect of the unlearned samples. In the Rectify stage, we utilize the induced graph of unlearned samples to approximate model gradients on the remaining dataset, which are then used to enhance the model performance on the remaining dataset.
Graph Unlearning via Parameter Masking
Previous research indicates that certain neurons or parameters are critical in memorizing specific samples (Maini et al. 2023), and that gradient manipulation can prevent the memorization of noisy information (Chen et al. 2021). Motivated by these findings, we propose to achieve efficient unlearning by directly masking the parameters responsible for memorizing the data to be unlearned, as described below:
$\widehat{\omega}_{r,j}=\begin{cases}0,&j\in M\\ \omega_{j}^{*},&j\notin M\\ \end{cases}$ (3)
where $\omega^{*}$ represents the optimal parameters on $D$, and $M$ denotes the set of parameters responsible for memorizing the unlearned dataset. To explore the effectiveness of the masking strategy, we provide theoretical justification for its ability to achieve graph unlearning as follows:
Theorem 1.
For a GNN model, if we approximate the FIM $F$ with its diagonal $diag(F)$, and assume all elements of $diag(F)$ are strictly positive, the mean squared distance between the optimal model parameters trained on $D_{r}$ and the parameters obtained from (3), denoted as $Q=\frac{1}{\left|\omega\right|}\sum_{j}{\left(\omega_{r,j}^{*}-\,\,\widehat{\omega}_{r,j}\right)}^{2}$, has the following upper bound:
$Q\leqslant\frac{1}{\left|\omega\right|}\left(c_{1}\sum_{j\in M}{\frac{1}{F_{D_{r},jj}^{2}}}+c_{2}\sum_{j\notin M}{\frac{F_{D_{f},jj}^{2}}{F_{D,jj}^{2}F_{D_{r},jj}^{2}}}+c_{3}\right)$ (4)
where $c_{1}$, $c_{2}$, $c_{3}$ are constants, and $|\omega|$ denotes the number of parameters.
The proof of Theorem 1 is provided in the Appendix. According to Theorem 1, adding the $j$-th parameter to $M$ will increase the upper bound of $Q$ by $\frac{1}{\left|\omega\right|}\left(\frac{c_{1}}{F_{D_{r},jj}^{2}}-\frac{c_{2}F_{D_{f},jj}^{2}}{F_{D,jj}^{2}F_{D_{r},jj}^{2}}\right)$. This increase is less than 0 when $\frac{F_{D_{f},jj}}{F_{D,jj}}>\sqrt{\frac{c_{1}}{c_{2}}}=c$, indicating that the parameter $w_{j}$ is much more important for $D_{f}$ than for $D$. Therefore, masking parameters that are crucial for unlearned samples but not for others can reduce the upper bound of $Q$, facilitating the forgetting of the unlearned dataset.
Erase: Neighborhood-Aware Parameter Editing
Despite the above analysis theoretically guaranteeing the effectiveness of the masking strategy in achieving unlearning, it has two major limitations First, the masking strategy overlook the extent to which $\frac{F_{D_{f},jj}}{F_{D,jj}}$ exceeds $c$. In particular, parameters for which $\frac{F_{D_{f},jj}}{F_{D,jj}}$ significantly exceeds $c$ are much more critical for the unlearned dataset compared to those where $\frac{F_{D_{f},jj}}{F_{D,jj}}$ only slightly exceeds $c$. Second, the masking strategy overlooks the impact of unlearned samples on their neighbors through message propagation.
To address the above limitations, we propose the Erase strategy as follows:
$\widehat{\omega}_{r,j}=\begin{cases}a\frac{F_{D,jj}}{F_{D_{f},jj}}\omega_{j}^{*},&F_{D_{f},jj}>\gamma F_{D,jj},\\ b\frac{{F_{D,jj}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}\omega_{j}^{*},&F_{D_{f},jj}F_{D_{k},jj}>{\eta F_{D,jj}}^{2}\\ &and\ F_{D_{f},jj}\leq\gamma F_{D,jj},\\ \omega_{j}^{*},&otherwise.\\ \end{cases}$ (5)
where $a$, $b$, $\gamma$, and $\eta$ are hyperparameters.
The advantages of the above design are threefold. First, it considers the impact of unlearned samples on their neighbors. According to the properties of the FIM, a large ratio of $\frac{F_{D_{f},jj}}{F_{D,jj}}$ or $\frac{F_{D_{k},jj}}{F_{D,jj}}$ indicates that the parameter is crucial for $D_{f}$ or $D_{k}$, but not for $D$. Drawing inspiration from Theorem 1, if a parameter is crucial for both $D_{f}$ and $D_{k}$, but not for other samples, i.e., satisfying $F_{D_{f},jj}F_{D_{k},jj}>\eta{F_{D,jj}}^{2}$, we consider that the parameter has been influenced by message propagation from $D_{f}$. In such cases, we modify these parameters to forget the influence of unlearned samples.
Second, when modifying parameters, we account for the extent to which they are influenced by unlearned samples. Specifically, rather than directly masking the parameters, we use the coefficients $\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\frac{{F_{D,jj}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$ to determine the degree of modification. This implies that a large $\frac{F_{D_{f},jj}}{F_{D,jj}}$ or $\frac{{F_{D_{f},jj}F_{D_{k},jj}}}{{F_{D,jj}}^{2}}$ necessitates substantial modifications to $\omega_{j}^{*}$ , while a small value requires only minor modifications.
Finally, we balance the two forgetting objectives: forgetting $D_{f}$ and forgetting the influence of $D_{f}$ on $D_{k}$. One of these objectives may dominate the other due to the potential difference in magnitude between $\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\frac{{F_{D,jj}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$. To mitigate this, we introduce two balancing coefficients $a$ and $b$ to balance these two forgetting objectives.
Note the choice of hyperparameters may vary depending on the dataset. To facilitate hyperparameter selection, we adaptively select them for different datasets. Specifically, $\gamma$ is set as the top $m$value of $\frac{F_{D_{f}}}{F_{D}}$, and $\eta$ is set as the top $m$value of $\frac{F_{D_{f}}F_{D_{k}}}{{F_{D}}^{2}}$. This ensures that $m$of the parameters satisfy $F_{D_{f},jj}>\gamma F_{D,jj}$ and $m$satisfy $F_{D_{f},jj}F_{D_{k},jj}>{\eta F_{D,jj}}^{2}$. Furthermore, we set $a=\gamma$ and $b=\eta$ to balance the magnitude between $\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\frac{{F_{D_{j}j}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$.
Algorithm 1 ETR: Erase then Rectify
1:Unlearned dataset $D_{f}$; Dataset $D_{k}$; GNN $f_{\mathcal{G}}$; Optimal parameters $\omega^{*}$; Gradient $\nabla_{\omega^{*}}L_{D}$; Hyperparameters $m$ and $\lambda$.
2:Erase
3:Calculate the gradients $\nabla_{\omega^{*}}L_{D_{f}}$ and $\nabla_{\omega^{*}}L_{D_{k}}$.
[…]
5:Obtain $\gamma=top\mbox{-}m\%(\frac{F_{D_{f}}}{F_{D}})$ and $\eta=top\mbox{-}m\%(\frac{F_{D_{f}}F_{D_{k}}}{{F_{D}}^{2}})$
6:for $j\ in\ range\ |\omega|$ do
7:     Obtain $\alpha_{j}=\frac{F_{D,jj}}{F_{D_{f},jj}}$ and $\beta_{i}=\frac{{F_{D,jj}}^{2}}{F_{D_{f},jj}F_{D_{k},jj}}$
8:     if $F_{D_{f},jj}>\gamma F_{D,jj}$ then
9:         $\widehat{\omega}_{r,j}=\alpha_{j}\gamma\omega_{j}^{*}$
10:     else if $F_{D_{f},jj}F_{D_{k},jj}>{\eta F_{D,jj}}^{2}$ then
11:         $\widehat{\omega}_{r,j}=\beta_{j}\eta\omega_{j}^{*}$
12:     else
13:         $\widehat{\omega}_{r,j}=\omega_{j}^{*}$
14:     end if
15:end for
16:Rectify
17:Calculate the gradient $\nabla_{\widehat{\omega}_{r}}L_{D_{k}}$.
[…]
20:$\omega^{\prime}$.
Backbone Method PubMed CiteSeer Cora CS Physics
F1 T F1 T F1 T F1 T F1 T
GCN Retrain 89.74±0.27 39.1 76.76±0.72 14.1 84.35±1.14 7.5 92.63±0.30 285.8 95.90±0.10 856.7
BEKM 80.03±0.72 68.2 61.11±2.27 12.7 66.42±3.33 11.0 89.03±0.49 71.7 94.53±0.19 152.7
BLPA 81.63±0.83 66.1 57.69±2.31 13.0 61.51±3.48 11.5 88.88±0.56 68.2 94.45±0.24 142.8
GIF 73.66±0.26 0.4 62.88±1.34 0.2 75.87±1.86 0.2 85.19±0.56 1.3 91.37±0.23 3.3
SR 88.04±0.25 29.0 75.35±0.70 8.6 81.25±0.95 8.6 91.77±0.28 49.0 94.72±0.14 166.4
Fast 88.11±0.24 29.1 75.83±0.88 8.8 80.70±1.70 8.4 91.64±0.29 47.8 94.74±0.16 138.3
MEGU 88.26±0.47 0.3 76.04±0.93 0.2 82.62±1.05 0.2 92.39±0.51 0.5 95.40±0.20 0.9
ETR 89.48±0.34 0.02 77.09±0.60 0.02 83.54±0.92 0.02 92.72±0.37 0.04 95.71±0.15 0.07
GAT Retrain 88.42±0.36 50.2 76.52±1.00 16.3 82.29±1.00 9.0 92.57±0.43 295.47 96.26±0.18 875.2
BEKM 70.30±0.77 102.2 49.64±1.36 23.5 56.79±4.83 20.7 81.10±1.39 114.3 92.04±0.43 214.5
BLPA 71.65±0.78 98.2 49.52±2.44 20.1 49.70±3.17 18.7 81.35±0.58 94.2 91.82±0.43 194.5
GIF 88.07±0.22 1.1 75.80±1.06 0.7 81.70±1.38 0.6 92.79±0.36 2.1 95.84±0.16 5.1
SR 86.67±0.41 33.6 75.92±0.88 12.0 76.05±1.78 12.3 89.49±0.64 54.8 94.76±0.41 149.3
Fast 86.60±0.33 35.4 75.89±0.93 12.4 75.46±0.93 12.3 89.70±0.95 55.3 94.86±0.45 145.5
MEGU 85.23±0.47 0.3 76.07±0.79 0.2 82.84±0.72 0.2 91.74±0.32 0.6 95.14±0.22 1.2
ETR 88.32±0.22 0.03 76.31±1.59 0.03 79.70±0.85 0.03 92.45±0.13 0.05 96.19±0.18 0.09
Table 1: Performance comparison in terms of F1 score and running time, where F1 indicates the F1 score, T indicates the running time (in seconds). The best experimental results are highlighted in bold, while the second-best results are underscored.
Rectify: Model Utility Enhancement
Although the Erase strategy mitigates the impact on the model’s predictive performance by considering the extent to which unlearned samples influence parameters, the parameter editing approach may still negatively affect the GNN’s performance on $D_{r}$ to some degree. In this part, we propose the Rectify strategy to enhance GNN performance on $D_{r}$ without requiring access to the entire training dataset, which minimizes computational overhead while ensuring data privacy and security.
Firstly, we estimate the gradients of the edited model on $D_{r}$ using the induced graph of $D_{f}$ Specifically, we assume that we can store the gradients from the model’s final iteration of the training phase. After removing $D_{f}$, the neighbor structure of $D_{r}\setminus D_{k}$ remains unchanged, while only the neighbor structures of $D_{f}$ and $D_{k}$ undergo changes. Additionally, during the Erase stage, the modified parameters are crucial only for $D_{f}$ and $D_{k}$, but not for $D_{r}\setminus D_{k}$. Therefore, for nodes in $D_{r}\setminus D_{k}$, their gradients can be considered nearly unchanged, i.e.,
$\nabla_{\widehat{\omega}_{r}}l_{j}\approx\nabla_{\omega^{*}}l_{j},\forall j\in D_{r}\setminus D_{k}$ (6)
where $l_{j}$ denotes the loss of the GNN on the $j$-th sample. Through the aforementioned approximation, we can easily calculate the gradient of the edited model on $D_{r}$ as follows:
$\nabla_{\widehat{\omega}_{r}}L_{D_{r}}=$ ${\frac{1}{\left|D_{r}\right|}}\sum_{j\in D_{r}}{\nabla_{\widehat{\omega}_{r}}l_{j}}$ (7)
$\approx$ ${\frac{1}{\left|D_{r}\right|}}\left(\sum_{j\in D_{r}\setminus D_{k}}{\nabla_{\omega^{*}}l_{j}}+\sum_{j\in D_{k}}{\nabla_{\widehat{\omega}_{r}}l_{j}}\right)$ (8)
where $L_{D}$ denotes the loss function of the GNN on dataset $D$. Further, leveraging the gradient of the GNN on D, we obtain $\sum_{j\in D_{r}\setminus D_{k}}{\nabla_{\omega^{*}}l_{j}}=\left|D\right|\nabla_{\omega^{*}}L_{D}-\sum_{j\in D_{f}}{\nabla_{\omega^{*}}l_{j}}-\sum_{j\in D_{k}}{\nabla_{\omega^{*}}l_{j}}$. Finally, we estimate the gradient of the edited model on $D_{r}$ as follows:
$\nabla_{\widehat{\omega}_{r}}L_{D_{r}}={\frac{1}{\left|D_{r}\right|}}($ $\left|D\right|\nabla_{\omega^{*}}L_{D}-\left|D_{f}\right|\nabla_{\omega^{*}}L_{D_{f}}-$ (9)
$\left|D_{k}\right|\nabla_{\omega^{*}}L_{D_{k}}+\left|D_{k}\right|\nabla_{\widehat{\omega}_{r}}L_{D_{k}})$
[…]
$\omega^{\prime}=\widehat{\omega}_{r}-\lambda\nabla_{\widehat{\omega}_{r}}L_{D_{r}}$ (10)
where $\lambda$ is the hyperparameter. The pseudo-code for the ETR approach is presented in Algorithm 1.
For edge and feature unlearning, we can forget the impact of the unlearned edges or features and enhance model performance in a similar manner. The pseudo-code for edge and feature unlearning is provided in the Appendix.
Complexity Analysis
In this section, we analyze the complexity of ETR. Calculating $\nabla_{\omega^{*}}L_{D_{f}}$ with a time complexity of $\mathcal{O}(|D_{f}|)$. Similarly, calculating $\nabla_{\omega^{*}}L_{D_{k}}$ and $\nabla_{\widehat{\omega}_{r}}L_{D_{k}}$ with a time complexity of $\mathcal{O}(|D_{k}|)$. Computing the FIM has a time complexity of $\mathcal{O}(|\omega|)$. The Erase strategy can be executed in parallel, with a time complexity of $\mathcal{O}(|\omega|)$. Thus, the overall time complexity of ETR is $\mathcal{O}(|D_{f}|+|D_{k}|+|\omega|)$. Regarding space complexity, the GNN parameters have a space complexity of $\mathcal{O}(|\omega|)$. The node features have a space complexity of $\mathcal{O}(|D_{f}|d+|D_{k}|d)$. The edges have a space complexity of $\mathcal{O}(|\mathcal{E}_{f}|+|\mathcal{E}_{k}|)$, where $|\mathcal{E}_{f}|$ and $|\mathcal{E}_{k}|$ denote the number of edges in $D_{f}$ and $D_{k}$, respectively. Computing the FIM has a space complexity of $\mathcal{O}(|\omega|)$. Therefore, the overall space complexity of ETR is $\mathcal{O}(|D_{f}|d+|D_{k}|d+|\omega|+|\mathcal{E}_{f}|+|\mathcal{E}_{k}|)$. Notably, $D_{f}$ and $D_{k}$ are much smaller than $D$, resulting in low computational overhead for large-scale graphs.
Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of ETR. The experiments aim to answer the following research questions: RQ1: How does ETR perform in terms of model utility, unlearning efficiency, and unlearning efficacy? RQ2: How does ETR perform on large-scale graphs? RQ3: How do hyperparameters influence ETR’s performance? RQ4: How do different strategies in ETR contribute to its effectiveness?
Experimental Setup
Datasets
We conduct experiments on PubMed, CiteSeer, Cora (Yang, Cohen, and Salakhutdinov 2016), CS, Physics (Shchur et al. 2018), ogbn-arxiv, and ogbn-products (Hu et al. 2020). The statistics and detailed descriptions of these datasets are provided in the Appendix.
Baselines
We compare ETR with various baselines, including Retrain, GraphEraser-BEKM, GraphEraser-BLPA (Chen et al. 2022), GIF (Wu et al. 2023a), GUIDE-SR, GUIDE-Fast (Wang, Huai, and Wang 2023), and MEGU (Li et al. 2024). Detailed descriptions of these baselines are provided in the Appendix.
Settings
Following GIF, we partition each graph into a training subgraph comprising training nodes and a test subgraph containing the remaining nodes. For PubMed, CiteSeer, Cora, CS, and Physics, we randomly allocate 90% of the nodes to the training set. For ogbn-arxiv and ogbn-products, we use the splits provided in (Hu et al. 2020). The unlearning ratio is set to 5% for all tasks. We use two-layer GCN and GAT models with a hidden state dimension of 256, training for 100 epochs. All experiments are conducted on a single Nvidia A40 GPU. Each experiment is run ten times, and we report the average value and standard deviation.
Evaluation of Model Utility (RQ1)
We evaluate the F1 score for node classification on the remaining dataset. The results for node unlearning are shown in Table 1, while the results for edge and feature unlearning are provided in the Appendix. It can be observed that ETR consistently performs best in most cases, which can be attributed to the Rectify strategy that enhances model performance on the remaining dataset. Additionally, BEKM and BLPA exhibit suboptimal performance, consistent with the results reported in their original papers, likely due to disruptions in graph structure resulting from graph partitioning. GUIDE mitigates this problem, leading to improved performance. Among the baselines, MEGU achieves the best performance, which can be attributed to its training of the model’s predictive objective during unlearning.
Backbone Method PubMed CiteSeer Cora CS Physics
GCN Retrain 1.57 1.52 1.3 6.13 14.62
BEKM 0.8 0.6 0.45 5.87 14.87
BLPA 0.45 0.48 0.38 1.44 2.89
GIF 1.7 1 0.61 11.18 37.63
SR 1.48 1.23 1.2 2.09 4.97
Fast 1.44 1.23 1.2 2.36 5.1
MEGU 0.68 0.52 0.41 2.07 4.52
ETR 1.18 1.21 1.17 1.61 2.43
GAT Retrain 1.77 1.52 1.3 6.17 15.28
BEKM 0.78 0.6 0.45 5.87 14.86
BLPA 0.48 0.48 0.4 1.6 3.53
GIF 2.88 0.8 0.72 5.89 14.3
SR 1.52 1.23 1.2 2.2 5.21
Fast 1.52 1.23 1.2 2.41 5.32
MEGU 0.94 0.57 0.45 3.62 7.64
ETR 1.2 1.21 1.17 1.54 2.3
Table 2: The memory overhead (GB) of different methods.
Backbone Method PubMed CiteSeer Cora CS Physics
GCN BEKM 5e-2 2e-3 8e-3 2e-2 2e-2
BLPA 5e-2 2e-3 8e-3 2e-2 1e-2
GIF 7e-2 4e-3 2e-2 3e-2 1e-2
SR 9e-2 8e-3 2e-2 8e-3 2e-3
Fast 8e-2 9e-3 2e-2 8e-3 2e-3
MEGU 1e-2 3e-3 3e-2 1e-3 7e-4
ETR 3e-2 5e-4 9e-4 4e-4 2e-4
GAT BEKM 9e-3 1e-2 2e-2 9e-3 5e-4
BLPA 9e-3 1e-2 2e-2 8e-3 5e-4
GIF 6e-2 7e-3 2e-2 2e-3 3e-2
SR 7e-2 1e-2 1e-2 4e-3 7e-2
Fast 7e-2 1e-2 1e-2 4e-3 7e-2
MEGU 9e-2 4e-2 3e-3 2e-2 1e-2
ETR 3e-2 1e-3 2e-3 4e-4 2e-4
Table 3: The root mean square distance between the parameters and those of Retrain.
Evaluation of Unlearning Efficiency (RQ1)
The runtime results are presented in Table 1. ETR consistently outperforms all baselines, which can be attributed to its training-free nature. ETR reduces runtime by thousands of times compared to Retrain, demonstrating its superior efficiency. While GraphEraser and GUIDE are also faster than Retrain in most cases, they still require considerable time for unlearning due to the need for retraining multiple models. Although GIF and MEGU improve runtime efficiency compared to other baselines, they still fall short of ETR. Specifically, ETR reduces runtime by 38 times compared to GIF and 12 times compared to MEGU.
The memory overhead results are shown in Table 2. ETR performs comparably to the baselines on small datasets and outperforms them on larger datasets (Photo and Computer). This is because memory overhead on small datasets is primarily driven by model parameters, limiting the advantages of ETR. However, ETR reduces memory overhead on large datasets by not requiring the entire training dataset, decreasing it by 3.4 times compared to Retrain. Among the baselines, BLPA performs best due to its use of graph partitioning, which reduces memory overhead but significantly degrades model utility. Overall, ETR achieves an excellent trade-off between unlearning efficiency and model utility.
Figure 2: Adversarial edge unlearning on the Cora dataset.
Bone Method ogbn-arxiv ogbn-products
F1 T S D F1 T S D
GCN Retrain 67.5 152 3.7 - 73.6 519 11.2 -
BEKM 64.8 1290 2.9 8e-1 OM OM OM OM
BLPA 63.9 1323 0.9 8e-1 OM OM OM OM
GIF 62.1 7 19.7 1 OM OM OM OM
MEGU 65.6 2 5.9 4e-2 71.3 7 26.8 6e-2
ETR 66.4 0.08 1.2 2e-2 73.8 0.5 1.2 9e-3
GAT Retrain 67.3 217.9 6 - 73.9 709 11.2 -
BEKM 65.3 1804 3.0 6e-1 OM OM OM OM
BLPA 64.7 1658 2.5 6e-1 OM OM OM OM
GIF 66.1 12 34.5 5e-1 OM OM OM OM
MEGU 58.5 3 11.2 4e-3 63.7 8 26.9 5e-2
ETR 67.8 0.1 1 2e-2 73.9 0.7 1.3 2e-2
Table 4: Performance comparison on large-scale graphs, where S denotes memory overhead, D indicates parameter distance, and “OM” represents “out of memory”.
Evaluation of Unlearning Efficacy (RQ1)
We assess unlearning efficacy using both direct and indirect evaluation methods. For direct evaluation, we measure the root mean square distance between the parameters of the unlearned model and those of the retrained model. As shown in Table 3, ETR consistently performs best in most cases, indicating that its parameters are closest to those obtained by retraining from scratch, thereby demonstrating ETR’s effectiveness in achieving graph unlearning.
For indirect evaluation, we follow the approach used by GIF and MEGU, adding adversarial edges between nodes of different categories in the training graph. These edges serve as unlearning targets, and we then assess the utility of the unlearned model. As shown in Figure 2, the performance of the vanilla trained model declines with increasing attack ratios, while GIF, MEGU, and ETR maintain or even improve performance. This demonstrates the effectiveness of these methods. Notably, ETR consistently outperforms the other baselines, highlighting its strong unlearning efficacy.
Evaluation on Large Scale Graphs (RQ2)
We conduct experiments on the ogbn-arxiv and ogbn-products datasets to assess ETR’s effectiveness and efficiency on large-scale graphs, with the results shown in Table 4. GUIDE encounters out-of-memory issues on both datasets, while GraphEraser and GIF also experience out-of-memory issues on the ogbn-products dataset. ETR demonstrates superior performance in terms of model utility, unlearning efficiency, and unlearning efficacy on large-scale graphs. Specifically, ETR shows an average utility improvement of 9.13% over MEGU, reduces runtime by thousands of times compared to Retrain, and decreases memory overhead by 7 times compared to Retrain and 15 times compared to MEGU. In terms of unlearning efficacy, ETR performs best in most cases. Overall, ETR achieves better performance and efficiency compared to other methods.
Figure 3: Performance with varying hyperparameters.
Hyperparameter Analysis (RQ3)
The hyperparameters $m$ and $\lambda$ are crucial for ETR’s performance. The value of $m$ determines how much the parameters are modified, thereby influencing the extent of model forgetting. The value of $\lambda$ controls the extent of model rectification. In this section, we investigate their impact on ETR’s performance using the PubMed, CiteSeer, and Cora datasets.
We investigate how different values of $m$ affect the performance of the Erase strategy, with results shown in Figure 3. It can be observed that increasing $m$ generally improves model performance, but performance declines if $m$ becomes too large. This is because larger $m$ values help the model forget unlearned samples effectively, while excessively large values can also lead to the loss of essential knowledge. Typically, $m$ values between 8‰ and 16‰ provide the best results for the Erase strategy. Notably, the model performs better compared to $m=0$, confirming that the Erase strategy effectively forgets unlearned samples and supports Theorem 1.
We tune $\lambda$ within the range of [0, 0.9] to investigate how it affects ETR’s performance, with results shown in Figure 3. It can be observed that ETR’s performance generally improves as $\lambda$ increases. However, performance declines if $\lambda$ becomes too large. This is because larger $\lambda$ values aid in model rectification, while excessively large values can also lead to excessive rectification. Optimal performance for ETR is typically achieved with $\lambda$ values between 0.3 and 0.5.
Method PubMed CiteSeer Cora CS Physics
w/ mask 88.91 77.00 82.51 92.44 95.80
w/o mes 89.30 76.91 82.44 92.53 95.70
ETR 89.48 77.09 83.54 92.72 95.71
Table 5: Performance comparison between model variants.
Metric PubMed CiteSeer Cora CS Physics
AD 5.7E-06 8.1E-07 1.7E-06 5.3E-06 1.6E-04
RD 0.23 0.09 0.07 0.36 0.22
Table 6: The difference between the approximate gradient and the true gradient.
Ablation Studies (RQ4)
In this section, we investigate the effectiveness of different strategies in ETR. The Erase is crucial for model forgetting, while the Rectify is critical for maintaining model utility. The results in Figure 3 indicate that increasing m improves performance compared to $m=0$, demonstrating the effectiveness of the Erase strategy in forgetting the unlearned samples. The results in Figure 3 show that increasing $\lambda$ also improves performance compared to $\lambda=0$, indicating that the Rectify strategy effectively enhances the model performance on the remaining dataset.
To further investigate the effectiveness of the Erase strategy, we compare ETR with two model variants, as shown in Table 5. Specifically, “w/ mask” refers to using the masking strategy, while “w/o mes” indicates not considering the impact of message propagation. The results show that ETR outperforms both variants in most cases, except for a slight decrease on the Physics dataset compared to “w/ mask”. This demonstrates that parameter editing is more effective than just applying a mask and emphasizes the importance of considering message propagation in graph unlearning.
In the Rectify stage, we use the induced graph of unlearned samples to approximate the gradient of the model on the remaining dataset. We measure the difference of this approximation using Absolute Difference (AD) and Relative Difference (RD), as shown in Table 6. The results show that both differences are minimal, indicating that the gradient approximation ensures both efficiency and accuracy.
Conclusion
In this paper, we investigate cost-effective graph unlearning and propose a training-free approach ETR to remove the influence of unlearned samples while maintaining the model performance on the remaining dataset. Specifically, in the Erase stage, we utilize the Fisher Information Matrix to identify parameters crucial for unlearned samples and their influence on other inter-connected samples. These parameters are then modified to forget the unlearned samples and their effects on other samples through message propagation. In the Rectify stage, we use the induced graph of the unlearned samples to efficiently approximate the model’s gradient on the remaining dataset. This gradient is then used to enhance the model performance on the remaining data. Extensive experiments on seven public datasets demonstrate that the proposed ETR method achieves an excellent trade-off between efficiency and effectiveness.
Appendix A Theoretical Analysis
Prior to proving Theorem 1, we first introduce some notation that will be used in the proof. Considering the architectural differences among various GNNs, for the sake of simplicity in our analysis, we consider the following general two-layer GNN architecture:
$H_{1}$ $=\sigma\left(PXW_{0}\right)$ (11)
$H_{2}$ $=PH_{1}W_{1}$ (12)
$Z$ $=softmax\left(H_{2}\right)$ (13)
where $P$ represents the message transformation matrix associated with the adjacency matrix, $\sigma(\cdot)$ denotes the activation function, $W_{0}$ and $W_{1}$ represent parameter matrices. In the node classification task, the loss function is given by:
$L=-{\frac{1}{\left|D\right|}}\sum_{i\in D}{\sum_{j=1}^{C}{Y_{i,j}\log\left(Z_{i,j}\right)}}.$ (14)
Based on the definition of the Fisher Information Matrix, the following properties are established:
$F_{D}$ $=\frac{1}{\left|D\right|}\sum_{i=1}^{\left|D\right|}{\nabla_{w}\log p\left(y_{i}|x_{i},w\right)\nabla_{w}\log p\left(y_{i}|x_{i},w\right)^{T}}$ (15)
$=\frac{1}{\left|D\right|}\sum_{i\in D_{f}}{\nabla_{w}\log p\left(y_{i}|x_{i},w\right)\nabla_{w}\log p\left(y_{i}|x_{i},w\right)^{T}}$ (16)
$+\frac{1}{\left|D\right|}\sum_{i\in D_{r}}{\nabla_{w}\log p\left(y_{j}|x_{j},w\right)\nabla_{w}\log p\left(y_{j}|x_{j},w\right)^{T}}$ (17)
$=\frac{\left|D_{f}\right|}{\left|D\right|}F_{D_{f}}+\frac{\left|D_{r}\right|}{\left|D\right|}F_{D_{r}}$ (18)
We denote the vectorization of the matrix $W$ as $vec(W)$ and define $\omega=vec(W_{0},W_{1})=concat(vec(W_{0}),vec(W_{1}))$, where $concat(\cdot,\cdot)$ denotes the concatenation of two vectors. We now present the proof of Theorem 1.
Proof of Theorem 1.
First, we compute the gradients of the backpropagation using the node classification loss function as follows:
$\frac{\partial L}{\partial W_{1}}$ $=H_{1}^{T}P^{T}\left(Z-Y\right)$ (19)
$\frac{\partial L}{\partial W_{0}}$ $=X^{T}P^{T}\left\{\left[\left(Z-Y\right)W_{1}^{T}P^{T}\right]\odot\sigma^{\prime}\left(PXW_{0}\right)\right\}$ (20)
According to the definition of the Fisher Information Matrix, we have: $F=vec\left(\frac{\partial L}{\partial W_{1}},\frac{\partial L}{\partial W_{0}}\right)vec\left(\frac{\partial L}{\partial W_{1}},\frac{\partial L}{\partial W_{0}}\right)^{T}$. For the sake of simplicity, we decompose $F$ by analyzing $W_{0}$ and $W_{1}$ separately. Define $F_{0}=vec\left(\frac{\partial L}{\partial W_{0}}\right)vec\left(\frac{\partial L}{\partial W_{0}}\right)^{T}$ and $F_{1}=vec\left(\frac{\partial L}{\partial W_{1}}\right)vec\left(\frac{\partial L}{\partial W_{1}}\right)^{T}$. Accordingly, we have: $F_{1}vec\left(W_{1}\right)=tr\left(H_{2}\left(Z-Y\right)\right)vec\left(\frac{\partial L}{\partial W_{1}}\right)$, and $F_{0}vec\left(W_{0}\right)=tr\left(W_{0}\frac{\partial L}{\partial W_{0}}\right)vec\left(\frac{\partial L}{\partial W_{0}}\right)$. We denote $b\coloneqq F\omega\approx\left[F_{0}W_{0};F_{1}W_{1}\right]$
According to Formulation (3), we have:
$\frac{1}{\left|\omega\right|}\sum_{j}{\left(\omega_{r,j}^{*}-\,\,\widehat{\omega}_{r,j}\right)}^{2}$ (21)
$=$ $\frac{1}{\left|\omega\right|}\left(\sum_{j\in M}{{\omega_{r,j}^{*}}^{2}}+\sum_{j\notin M}{\left(\omega_{r,j}^{*}-\,\,\omega_{j}^{*}\right)^{2}}\right)$ (22)
Using $diag(F)$ to approximate the Fisher information matrix $F$, the first term of Formulation (22) is given by:
$\sum_{j\in M}{{\omega_{r,j}^{*}}^{2}}=\sum_{j\in M}{\frac{b_{r,j}^{2}}{F_{D_{r},jj}^{2}}}\leqslant c_{1}\sum_{j\in M}{\frac{1}{F_{D_{r},jj}^{2}}}$ (23)
[…]
$\sum_{j\notin M}{\left(\omega_{r,j}^{*}-\,\,\omega_{j}^{*}\right)^{2}}$ (24)
$=$ $\sum_{j\notin M}{\left(\frac{b_{j}}{F_{D,jj}}-\,\,\frac{b_{r,j}}{F_{D_{r},jj}}\right)^{2}}$ (25)
$=$ $\sum_{j\notin M}{\left(\frac{b_{j}F_{D_{r},jj}-b_{r,j}F_{D,jj}}{F_{D,jj}F_{D_{r},jj}}\right)^{2}}$ (26)
$=$ $\sum_{j\notin M}{\left(\frac{b_{j}F_{D_{r},jj}-b_{r,j}\frac{\left|D_{f}\right|}{\left|D\right|}F_{D_{f},jj}-b_{r,j}\frac{\left|D_{r}\right|}{\left|D\right|}F_{D_{r},jj}}{F_{D,jj}F_{D_{r},jj}}\right)^{2}}$ (27)
$=$ $\sum_{j\notin M}{\left(\left(b_{j}-b_{r,j}\frac{\left|D_{r}\right|}{\left|D\right|}\right)\frac{1}{F_{D,jj}}-b_{r,j}\frac{\left|D_{f}\right|}{\left|D\right|}\frac{F_{D_{f},jj}}{F_{D,jj}F_{D_{r},jj}}\right)^{2}}$ (28)
$\leqslant$ $\sum_{j\notin M}{\left(b_{j}-b_{r,j}\frac{\left|D_{r}\right|}{\left|D\right|}\right)^{2}}\frac{2}{F_{D,jj}^{2}}$ (29)
$+$ $\sum_{j\notin M}{\left(b_{r,j}\frac{\left|D_{f}\right|}{\left|D\right|}\right)^{2}}\frac{F_{D_{f},jj}^{2}}{F_{D,jj}^{2}F_{D_{r},jj}^{2}}$ (30)
$\leqslant$ $c_{2}\sum_{j\notin M}{\frac{F_{D_{f},jj}^{2}}{F_{j,j}^{2}F_{r,jj}^{2}}}+c_{3}$ (31)
[…]
Therefore, we can obtain:
$\frac{1}{\left|\omega\right|}\sum_{j}{\left(\omega_{r,j}^{*}-\,\,\widehat{\omega}_{r,j}\right)}^{2}$ (32)
$=$ $\frac{1}{\left|\omega\right|}\left(\sum_{j\in M}{{\omega_{r,j}^{*}}^{2}}+\sum_{j\notin M}{\left(\omega_{r,j}^{*}-\,\,\omega_{j}^{*}\right)^{2}}\right)$ (33)
$\leqslant$ $\frac{1}{\left|\omega\right|}\left(c_{1}\sum_{j\in M}{\frac{1}{F_{D_{r},jj}^{2}}}+c_{2}\sum_{j\notin M}{\frac{F_{D_{f},jj}^{2}}{F_{D,jj}^{2}F_{D_{r},jj}^{2}}}+c_{3}\right)$ (34)
Appendix B Datasets
Dataset Statistics
Dataset #Nodes #Edges #Features #Classes
PubMed 19,717 88,651 500 3
CiteSeer 3,327 9,228 3,703 6
Cora 2,708 10,556 1,433 7
CS 18,333 163,788 6,805 15
Physics 34,493 495,924 8,415 5
ogbn-arxiv 169,343 1,166,243 128 40
ogbn-products 2,449,029 61,859,140 100 47
Table 7: Statistics of datasets.
Dataset Descriptions
  • PubMed, CiteSeer, and Cora. PubMed, CiteSeer, and Cora are three standard citation network benchmark datasets. In these datasets, the nodes correspond to research papers, and the edges represent the citations between papers. The node features are derived from the bag-of-words representation of the papers, while the node labels indicate the academic topics of the papers.
  • CS and Physics. CS and Physics are two co-authorship networks focused on research papers. In these datasets, nodes represent authors of research papers, while edges denote the co-authorship relationships among them. Node features derive from the keywords associated with each author’s publications, and node labels indicate the most active fields of study for each author.
  • ogbn-arxiv. ogbn-arxiv is a large-scale citation network comprising Computer Science papers from arXiv. In this dataset, nodes represent individual arXiv papers, while edges signify the citation relationships between them. Node features are generated by averaging the embeddings of the words found in the title and abstract of each paper. Node labels denote the subject areas of the papers.
  • ogbn-products. ogbn-products is a large-scale co-purchasing network of Amazon products. In this dataset, nodes represent individual products sold on Amazon, while edges indicate pairs of products that were purchased together. Node features are derived by extracting bag-of-words features from product descriptions. Node labels denote the product categories.
Appendix C Baselines
  • Retrain: Retrain is an exact graph unlearning method that entails retraining the model from scratch on the remaining dataset upon receiving an unlearning request.
  • GraphEraser: GraphEraser is the first approximate graph unlearning method, adhering to the SISA paradigm. It partitions the training graph into multiple shards, trains a model on each shard, and subsequently aggregates the various submodels. Upon receiving an unlearning request, only the affected submodels need to be retrained. GraphEraser-BEKM and GraphEraser-BLPA denote the application of different graph partitioning methods.
  • GIF: GIF utilizes influence functions to approximate the parameter changes resulting from an unlearning request. It subsequently updates the model parameters based on these estimated changes to facilitate graph unlearning.
  • GUIDE: GUIDE is specifically designed for inductive graph unlearning and follows the SISA paradigm. GUIDE takes into account fairness and balance in graph partitioning and addresses the disruptions caused to the graph structure. GUIDE-Fast and GUIDE-SR represent the application of different graph partitioning methods.
  • MEGU: MEGU integrates the model’s predictive capability and forgetting ability within a unified optimization framework, thereby facilitating mutual benefits between these two objectives.
Algorithm 2 ETR for Edge and Feature Unlearning
1:Influenced dataset $D_{i}$; GNN $f_{\mathcal{G}}$; Optimal parameters $\omega^{*}$; Gradient $\nabla_{\omega^{*}}L_{D}$; Hyperparameters $m$ and $\lambda$.
2:Erase
3:Calculate the gradients $\nabla_{\omega^{*}}L_{D_{i}}$.
[…]
5:Obtain $\gamma=top\mbox{-}k\%(\frac{F_{D_{i}}}{F_{D}})$.
6:for $j\ in\ range\ |\omega|$ do
7:     Obtain $\alpha_{j}=\frac{F_{D,jj}}{F_{D_{i},jj}}$
8:     if $F_{D_{i},jj}>\gamma F_{D,jj}$ then
9:         $\widehat{\omega}_{r,j}=\alpha_{j}\gamma\omega_{j}^{*}$
10:     else
11:         $\widehat{\omega}_{r,j}=\omega_{j}^{*}$
12:     end if
13:end for
14:Rectify
15:Calculate the gradient $\nabla_{\widehat{\omega}_{r}}L_{D_{k}}$.
[…]
18:$\omega^{\prime}$.
Appendix D ETR for Edge Unlearning and Feature Unlearning
For edge and feature unlearning, it is necessary to forget the impact of the unlearned edges or features on other nodes. We denote the subgraph affected by the unlearned edges or features as $D_{i}$. Consequently, we modify the parameters that are crucial for $D_{i}$ but not for the remaining dataset, as detailed in Algorithm 2. Additionally, we approximate the gradient of the model on the remaining dataset as follows:
$\nabla_{\widehat{\omega}_{r}}L_{D_{r}}={\frac{1}{\left|D_{r}\right|}}($ $\left|D\right|\nabla_{\omega^{*}}L_{D}-\left|D_{i}\right|\nabla_{\omega^{*}}L_{D_{i}}+$ (35)
$\left|D_{i}\right|\nabla_{\widehat{\omega}_{r}}L_{D_{i}})$ (36)
Furthermore, we utilize this gradient to enhance the model performance on the remaining dataset, as demonstrated in Algorithm 2.
Appendix E Additional Experimental Results
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN Retrain 89.95±0.18 78.17±0.84 86.64±0.54 93.48±0.23 96.45±0.18 68.07±1.38
BEKM 78.90±0.48 59.85±1.70 65.68±3.01 87.62±0.64 94.66±0.37 64.90±0.25
BLPA 80.52±0.44 56.52±2.86 60.66±2.90 87.48±0.74 94.46±0.34 63.60±0.56
GIF 83.20±0.22 72.55±1.06 82.62±1.38 92.62±0.36 95.61±0.16 65.76±1.06
SR 87.36±0.28 78.26±0.79 84.46±0.93 92.14±0.14 94.91±0.12 OOM
Fast 87.36±0.24 78.44±0.87 84.65±0.89 92.14±0.17 94.95±0.12 OOM
MEGU 87.81±0.23 77.27±0.81 86.53±0.80 93.23±0.40 96.01±0.11 66.97±1.67
ETR 89.88±0.21 79.28±0.76 87.23±0.76 93.48±0.26 96.43±0.16 68.10±0.63
GAT Retrain 88.60±0.26 77.66±0.91 84.13±1.44 93.08±0.17 96.65±0.11 69.03±0.60
BEKM 70.01±0.91 53.48±2.34 48.93±2.70 82.19±0.47 91.09±0.44 64.79±0.22
BLPA 71.26±1.17 51.59±2.86 49.89±6.71 81.80±0.62 90.83±0.27 64.37±0.26
GIF 88.04±0.33 76.91±1.35 85.06±1.18 93.59±0.22 96.26±0.19 67.49±1.46
SR 85.74±0.49 76.25±0.85 79.82±1.54 90.51±0.68 94.46±0.98 OOM
Fast 85.81±0.64 75.65±0.86 79.45±2.01 90.60±0.76 93.81±1.31 OOM
MEGU 84.24±0.40 75.44±0.88 86.61±1.11 92.32±0.31 95.09±0.19 59.69±1.69
ETR 88.21±0.48 77.87±0.93 84.32±0.91 93.40±0.27 96.61±0.17 69.40±0.46
Table 8: Performance comparison in terms of model utility in the edge unlearning task.
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN Retrain 89.96±0.26 78.35±0.62 87.16±0.42 93.59±0.13 96.55±0.11 68.59±1.27
BEKM 79.94±0.50 58.26±3.50 60.89±6.51 89.22±0.52 94.67±0.17 65.01±0.24
BLPA 81.60±0.59 54.17±3.23 61.33±2.85 88.79±0.33 94.58±0.24 63.87±0.33
GIF 84.60±0.26 73.54±0.48 84.24±1.00 93.06±0.27 95.63±0.13 65.52±1.30
SR 87.44±0.24 77.03±0.50 78.19±4.21 91.19±0.36 95.00±0.09 OOM
Fast 87.44±0.35 76.97±0.76 78.23±3.76 91.32±0.21 95.00±0.10 OOM
MEGU 87.87±0.17 77.60±0.51 85.65±0.71 93.50±0.17 96.01±0.14 66.70±1.95
EAR 90.17±0.24 78.98±0.71 87.08±0.43 93.63±0.22 96.54±0.11 68.27±0.92
GAT Retrain 88.52±0.26 77.66±0.79 85.20±0.81 93.25±0.46 96.57±0.20 68.76±0.77
BEKM 69.31±1.15 52.76±2.92 59.45±3.08 82.52±0.70 91.15±0.41 65.38±0.23
BLPA 70.84±0.95 48.32±1.71 55.17±3.94 82.72±0.38 91.17±0.36 64.82±0.25
GIF 88.11±0.37 77.33±1.30 85.35±1.43 93.04±0.22 96.58±0.18 64.38±11.07
SR 85.65±0.30 75.35±0.87 73.91±3.47 82.55±3.33 91.73±1.89 OOM
Fast 85.34±0.28 75.83±0.73 73.25±4.07 82.57±4.09 90.36±2.70 OOM
MEGU 84.20±0.48 75.59±0.91 87.31±1.06 92.24±0.24 95.50±0.14 59.40±1.75
EAR 88.45±0.41 77.84±0.96 84.58±0.93 93.22±0.25 96.66±0.15 69.54±0.35
Table 9: Performance comparison in terms of model utility in the feature unlearning task.
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN Retrain 42.82s 15.64s 8.31s 357.49s 1057.06s 173.59s
BEKM 76.60s 12.92s 10.92s 66.88s 138.76s 1314.22s
BLPA 65.70s 13.16s 11.36s 69.25s 140.64s 1424.94s
GIF 0.40s 0.22s 0.21s 1.30s 1.92s 6.87s
SR 30.81s 8.71s 8.80s 52.74s 149.74s OOM
Fast 30.71s 8.65s 8.71s 52.15s 152.54s OOM
MEGU 0.30s 0.20s 0.19s 0.53s 1.04s 2.77s
ETR 0.01s 0.01s 0.01s 0.03s 0.05s 0.06s
GAT Retrain 54.58s 18.28s 9.96s 369.58s 1079.54s 238.50s
BEKM 112.55s 20.24s 17.11s 94.77s 182.37s 1435.04s
BLPA 96.90s 20.47s 17.72s 97.58s 188.74s 1558.15s
GIF 1.11s 0.68s 0.64s 2.10s 5.10s 11.56s
SR 35.80s 12.97s 12.53s 56.25s 158.29s OOM
Fast 35.54s 12.22s 12.24s 58.30s 157.83s OOM
MEGU 0.37s 0.23s 0.22s 0.66s 1.36s 3.64s
ETR 0.02s 0.02s 0.02s 0.04s 0.06s 0.08s
Table 10: Performance comparison in terms of unlearning efficiency in the edge unlearning task.
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN Retrain 42.22s 15.60s 7.88s 322.21s 943.82s 168.91s
BEKM 71.08s 12.95s 11.52s 71.71s 142.22s 1327.08s
BLPA 68.97s 13.67s 11.63s 69.52s 143.42s 1320.60s
GIF 0.39s 0.22s 0.18s 1.30s 1.94s 6.88s
SR 30.34s 8.44s 8.73s 52.22s 148.24s OOM
Fast 30.92s 8.62s 8.35s 52.06s 147.64s OOM
MEGU 0.31s 0.20s 0.19s 0.53s 1.05s 2.77s
ETR 0.01s 0.01s 0.01s 0.03s 0.05s 0.06s
GAT Retrain 53.17s 17.84s 9.57s 332.11s 963.44s 235.38s
BEKM 105.58s 20.42s 18.14s 100.69s 198.04s 1610.27s
BLPA 101.28s 21.49s 19.09s 108.43s 231.09s 1595.64s
GIF 1.12s 0.65s 0.59s 2.10s 5.09s 11.58s
SR 36.00s 12.73s 12.12s 57.75s 169.89s OOM
Fast 35.85s 12.25s 12.08s 55.81s 159.12s OOM
MEGU 0.37s 0.24s 0.22s 0.66s 1.37s 3.67s
ETR 0.02s 0.02s 0.02s 0.03s 0.06s 0.08s
Table 11: Performance comparison in terms of unlearning efficiency in the feature unlearning task.
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN BEKM 0.095 0.10 0.0098 0.43 0.049 1.48
BLPA 0.13 0.095 0.0088 0.43 0.050 1.53
GIF 0.072 0.0045 0.029 0.029 0.22 1.44
SR 0.085 0.051 0.016 0.0088 0.0019 OOM
Fast 0.083 0.050 0.016 0.0088 0.0019 OOM
MEGU 0.0085 0.0023 0.0033 0.0012 0.00059 0.041
ETR 0.015 0.00015 0.00021 0.00026 0.00011 0.013
GAT BEKM 0.010 0.015 0.030 0.015 0.0065 0.12
BLPA 0.011 0.015 0.031 0.015 0.005 0.12
GIF 0.10 0.0069 0.014 0.0017 0.0079 1.86
SR 0.036 0.012 0.011 0.031 0.086 OOM
Fast 0.034 0.014 0.011 0.031 0.086 OOM
MEGU 0.082 0.0012 0.0023 0.023 0.0022 0.0033
ETR 0.011 0.00040 0.00041 0.00029 0.00014 0.016
Table 12: Performance comparison in terms of unlearning efficacy in the edge unlearning task.
Backbone Method PubMed CiteSeer Cora CS Physics ogbn-arxiv
GCN BEKM 0.091 0.018 0.0092 0.035 0.055 1.43
BLPA 0.097 0.014 0.0092 0.036 0.059 1.52
GIF 0.072 0.0045 0.024 0.0098 0.035 1.55
SR 0.082 0.0080 0.0090 0.0087 0.0023 OOM
Fast 0.081 0.0079 0.0093 0.0088 0.0023 OOM
MEGU 0.0084 0.0024 0.019 0.00035 0.00058 0.041
ETR 0.0090 0.00011 0.00011 0.00023 0.00010 0.013
GAT BEKM 0.011 0.014 0.051 0.0051 0.0084 0.65
BLPA 0.011 0.013 0.055 0.0051 0.0069 0.66
GIF 0.10 0.0070 0.040 0.033 0.0089 1.79
SR 0.024 0.0036 0.0073 0.0042 0.097 OOM
Fast 0.025 0.0036 0.0072 0.0043 0.097 OOM
MEGU 0.081 0.0012 0.0023 0.023 0.012 0.0032
ETR 0.0061 0.00029 0.00023 0.00028 0.00014 0.012
Table 13: Performance comparison in terms of unlearning efficacy in the feature unlearning task.